OpenStudy (anonymous):

2 people walk 4 miles toward each other, one at a 3 mph and the other at 2mph. What is an equation that shows how long it will take before they meet?

OpenStudy (anonymous):

Distance = speed*time Distance1 = 3*time Distance2 = 2*time Distance1 + distance2 = 4 3*t + 2*t = 4 5t = 4 t = 4/5 They will meet at 4/5 of an hour which is 48 minutes. A would of travelled (4/5*3) 2.4 miles. B would of travelled (4/5*2) 1.6 miles. Check: 2.4+1.6 = 4 miles. The equation is: 3*t + 2*t = 4 simplified as 5t = 4.